# Entanglement engineering of optomechanical systems by reinforcement learning

### v2

@@ Introduction
At the present, a major limitation/challenge in entanglement engineering is the experimental observation design. Existing machine-learning based works use the full fidelity, i.e., the overlap between the current and target quantum states, as the observation metric. Applications range from the generation of two [22] and multi-qubit entangled states [23, 24] to specific many-body states [25, 26, 27] and single-particle quantum state engineering via deep reinforcement learning (RL) [28, 29]. However, full fidelity observation is not universally applicable in experiments. Moreover, obtaining the relationship between the entanglement and experimental observables is difficult. So far there have been no systematical methods to extract quantitative entanglement from experimental observation for arbitrary quantum systems [30, 31, 32], in spite of some initial exploration for specific systems. For example, an entanglement criterion for non-Gaussian states in coupled harmonic oscillators was developed [30]. Under the strong laser approximation, a Bell inequality was tested with photon counting [31], and stationary entanglement for Gaussian states was inferred from the continuous measurement of light only [32].
@@
In this paper, using quantum optomechanical systems with linear or nonlinear photon-phonon interactions as a paradigm, we develop a deep RL approach to entanglement engineering. For quantum control of optomechanical systems, most existing theoretical studies focused on Gaussian states or the linear interaction regime [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 38, 44, 45], with the primary goal of generating entanglement as quickly as possible (entanglement enhancement) [43, 38, 44, 45]. Previous control methods are mostly model-based: prior information about the system model is needed, such as the pulse method [33, 34, 35, 36, 37], time-continuous laser-driven approaches [38, 39], periodic modulations [40, 41, 42], optimal pulse protocols [43], linear quadratic-Gaussian (LQG) methods [38], and coherent feedback methods using auxiliary optical components [44, 45]. We note that there were two previous works [46, 47] on model-free RL for controlling and stabilizing a quantum system with an inverted harmonic potential and a double-well nonlinear potential, respectively, to a target state using weak-current measurements (WCMs) and partial state observation. However, these two works did not address entanglement control, while our work is developing a model-free deep-RL method to realize non-Gaussian entanglement engineering using only photon number counting from WCMs. To our knowledge, prior to our work, model-free deep RL feedback control to create and stabilize the entanglement with WCM observations had not been available.
@@
Results
Experimental proposal for entanglement engineering
@@ Results
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the cavity and mechanical mode, respectively, $\hat{a}^{\dagger}$ and $\hat{b}^{\dagger}$ are the corresponding creation operators. The frequency detuning of the cavity is $\Delta\equiv\omega_{L}-\omega_{c}$. The nonlinear coupling $g_{0}$ arises from the radiation pressure force between the light and the movable mirror (details given in Supplementary Note 2), and $\alpha_{L}$ is the real amplitude of the driven electromagnetic field. We set $g_{0}>\kappa$ so that the single-photon optomechanical coupling rate $g_{0}$ exceeds the coupling strength $\kappa$ between the cavity and the vacuum bath. This condition guarantees observable nonlinear quantum effects [61]. Under the strong laser approximation: $|\bar{\alpha}_{c}|\gg 1$, where $|\bar{\alpha}_{c}|$ is the amplitude of the light field inside the cavity induced by the strong laser, we have $\hat{a}\approx\bar{\alpha}_{c}+\delta\hat{a}$ with $\delta\hat{a}$ denoting the excitation or the shifted oscillator on top of the large coherent state with the amplitude $\bar{\alpha}_{c}$. The resulting linearized beam-splitter or “swap” Hamiltonian  [62, 55] is
@@
The open dissipative quantum optomechanics under the WCM obey the stochastic master equation (SME) (see Methods). The number $n_{\textnormal{traj}}$ of trajectories simulated from SME can be selected according to the following considerations. If the observable is some expected physical quantity, using one trajectory is sufficient to extract the information about the quantum state: $n_{\textnormal{traj}}=1$. Experimentally, WCMs are performed, encoding the Wiener process in the observation and resulting in a large variance from the expectation value. To reduce the variance, more quantum trajectories should be used. To make computations feasible, we use five trajectories: $n_{\textnormal{traj}}=5$.
@@
RL in linear quantum optomechanics
@@
with the associated expected photon number $\langle\hat{n}^{\textnormal{target}}_{p}\rangle=\langle\hat{P}_{1}\rangle=0.5$. Consequently, the reward function can be set as $R_{t}\equiv-|O_{t}-0.5|$, regardless of whether the observation $O_{t}$ is $\langle\hat{n}_{p}\rangle(t)$ or $\mathcal{I}(t)$. Because of the relatively small target value of the expected photon number: $\langle\hat{P}_{1}\rangle=0.5$, the variance $\mathcal{I}(t)$ in the WCM photocurrent can be reduced by a Gaussian filter [86] with the weak measurement rate $\eta\leq 1$. The Gaussian kernel parameters of the filter such as the filter interval and the variance can be numerically chosen to reduce the standard deviation of the measurement photocurrent into a certain range, e.g., about ten times larger than the mean value (details in Supplementary Note 6). The PPO agent applies an updated stochastic policy to the quantum optomechanical environment to maximize the accumulated reward, where the action $G(t)$ is proportional to the amplitude of the cavity mode: $G(t)=g_{0}\bar{\alpha}_{c}(t)$. The action can be controlled by an incident laser [63] and is continuous in a certain range, e.g., $G\in[-5,5]\omega_{m}$. The decay rate of cavity and mechanical modes $\kappa=0.01\,\omega_{m}$, $\gamma=0.01\,\kappa$, respectively, because the quality of the mechanical oscillator mode is generally better than that of the optical cavity or microwave resonator mode [61, 59, 58, 65, 66, 67].
@@
Our deep RL, a model-free learning method, is implemented in the measurement-based feedback control framework for entanglement engineering in open quantum optomechanics. Details about the PPO algorithm applied in the linear quantum optomechanics are presented in Supplementary Note 4. To appreciate its performance, we employ two benchmark methods for comparison: Bayesian [54, 47] and random control. Bayesian control [54, 47] is a state-based feedback control of the stochastic process as governed by the SME. In our case, the control law is given by $G(t)=-\lambda|\langle\hat{n}_{p}\rangle(t)-0.5|\omega_{m}$ with $\langle\hat{n}_{p}\rangle(t)$ being the observation, where the hyperparameter $\lambda$ can be numerically optimized based on the performance. If the observation is $\mathcal{I}(t)$, the control flow will be in the form $G(t)=-\lambda|\mathcal{I}(t)-0.5|\omega_{m}$, in which the Wiener process blocks the performance to some degree. In Bayesian control, the smaller the variance in the measured photocurrent, the better the performance. For the random control method, the flow is generated by a uniform distribution in the action range $G\in[-5,5]\,\omega_{m}$. Note that the actions $G$ of random control and deep RL are in the same range while the one of Bayesian control is determined by the hyperparameter $\lambda$ and the state-based observation value or the WCM photocurrent. To make a fair comparison, $\lambda_{opt}$ is optimized within the action range $G\in[-5,5]\omega_{m}$. Specifically, the optimized hyperparameter $\lambda_{opt}$ corresponds to the best performance of Bayesian control in the set $\lambda\in\{1,2,\ldots,\lambda_{max}\}$, where $\lambda_{max}$ is the maximum integer of $\lambda$ to guarantee the action range $G\in[-5,5]\,\omega_{m}$.
@@
RL in nonlinear quantum optomechanics
@@
Physical understanding of entanglement engineering through model-free deep RL
@@
Linear interaction regime
@@
Nonlinear interaction regime
@@
Weak continuous measurement
In an open quantum system, under WCM and quantum dissipation into the vacuum bath as well, a Wiener process occurs in the observable. More specifically, the Wiener process arises from the Gaussian-weighted projection over the eigenstates, which weakly extracts the partial information from the quantum system and induces stochastic disturbances in both the dynamical equation and observation. Such disturbances can avoid a complete quantum state collapse and provide the capability to extract the quantum information continuously in the time domain. However, the nonlinear stochastic process occurs in both quantum dynamical trajectories and the measurement photocurrent, making it challenging to control the quantum system through WCM continuously. (Backgrounds about WCM, deep RL, and quantum control are presented in Supplementary Note 1.)
@@ Discussion
Methods
Stochastic Master Equation
@@ Methods
where the Hamiltonian is $\tilde{H}=\tilde{H}_{bs}$ or $\tilde{H}_{nl}$ and $\rho$ is a density operator in the Hilbert space. Under the Born-Markov approximation [95, 96], which requires the system-bath coupling to be weak and the correlation time of the bath to be much shorter than a characteristic timescale of system-bath interactions, the Markovian master equation, i.e., the first two terms in the right-hand side of Eq. (Stochastic Master Equation), has the Lindblad form [95]. At absolute zero temperature, the following environmental operator $\mathcal{L}_{env}\;\rho$ can be introduced to describe the coupling between the system and vacuum bath: $\mathcal{L}_{env}\;\rho=\kappa\mathcal{D}(\hat{a})\rho+\gamma\mathcal{D}(\hat{b})\rho$, where the cavity and oscillator modes are coupled to the vacuum bath with the strength $\kappa$ and $\gamma$, respectively [56]. The deep RL results in the Lindblad master equation with the nonlinear interaction are presented in Supplementary Note 5.
The WCM process described by the last two terms in the right-hand side of Eq. (Stochastic Master Equation) is nonlinear and Markovian in the unconditional master equation [51] in $\rho$. Under WCM, a Wiener process $dW$ with a Gaussian distribution [51] arises from the Gaussian-weighted projection over the eigenstates that allows the quantum information to be extracted continuously in the time domain, subject to stochastic disturbances in the last term of Eq. (Stochastic Master Equation) and quantum decoherence in the penultimate term of Eq. (Stochastic Master Equation). (Supplementary Note 3 in SI provides a detailed derivation of the SME.) The Lindblad operator $\mathcal{D}$ and the measurement superoperator $\mathcal{H}$ in Eq. (Stochastic Master Equation) are given by
@@
Implementation details of deep RL
For simulating the linear or nonlinear quantum optomechanical system described by Eq. (Stochastic Master Equation), we use the “taylor1.5” solver from the SME solver in the QuTip’s package [97] with the tolerance $\textnormal{tol}=10^{-6}$ and time step size $dt=0.01\,\omega_{m}^{-1}$. The measurement current is simulated with the “homodyne” method, and the custom environment is constructed by the open-source platform OpenAI-Gym [98]. For RL simulations, we construct the PPO agent [48] by “stable-baselines3” [99] in the A2C [100] settings, where stochastic policy (actor) and the value function (critic) are modeled by two independent neural network function approximators, i.e., a set of fully connected feed-forward networks of dimensions $256\times 128\times 64$ and the hyperbolic tangent nonlinear activation function for each hidden layer. For the nonlinear quantum optomechanical configuration, in the target-utilization phase, the recurrent PPO agent outperforms the PPO agent, where both independent critic and actor networks are MLPs followed by one independent layer of LSTM with $256\times 128\times 64$ fully connected networks and $256$ hidden states. More details are described in Supplementary Note 6.
Data availability
The data generated in this study about training results can be found in this Zenodo repository: https://doi.org/10.5281/zenodo.12584159 [101].
Code availability
The codes used in this paper can be found in the repository:
https://github.com/liliyequantum/Entanglement-engineering-by-RL [102].
Supplementary information
Supplementary information for this study is available on GitHub:
https://github.com/liliyequantum/Entanglement-engineering-by-RL [102].
Acknowledgment
We thank Dr. Kanu Sinha for discussions and comments. This work was supported by AFOSR under Grants Nos. FA9550-21-1-0438 and FA9550-21-1-0186. The Quantum Collaborative, led by Arizona State University, also provided valuable expertise and resources for this work through seed funding. The Quantum Collaborative connects top scientific programs, initiatives, and facilities with prominent industry partners to advance the science and engineering of quantum information science.

### v3

@@ I Introduction
At the present, a major limitation/challenge in entanglement engineering is the experimental observation design. Existing machine-learning based works use the full fidelity, i.e., the overlap between the current and target quantum states, as the observation metric. Applications range from the generation of two [22] and multi-qubit entangled states [23, 24] to specific many-body states [25, 26, 27] and single-particle quantum state engineering via deep reinforcement learning (RL) [28, 29]. However, full fidelity observation is not universally applicable in experiments. Moreover, obtaining the relationship between the entanglement and experimental observables is difficult. So far there have been no systematical methods to extract quantitative entanglement from experimental observation for arbitrary quantum systems [30, 31, 32], in spite of some initial exploration for specific systems. For example, an entanglement criterion for non-Gaussian states in coupled harmonic oscillators was developed [30]. Under the strong laser approximation, a Bell inequality was tested with photon counting [31], and stationary entanglement for Gaussian states was inferred from the continuous measurement of light only [32]. Recently, conditional state tomography [33] offers a method to reconstruct the density matrix from the weak measurement current. However, it requires the use of a specific Hamiltonian form, and the finite time needed for the conditional density matrix to converge to the actual density matrix. This highlights that quantum state tomography in a model-free manner for measurement-based feedback control remains a significant challenge.
@@
In this paper, using quantum optomechanical systems with linear or nonlinear photon-phonon interactions as a paradigm, we develop a deep RL approach to entanglement engineering. For quantum control of optomechanical systems, most existing theoretical studies focused on Gaussian states or the linear interaction regime [34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 39, 45, 46], with the primary goal of generating entanglement as quickly as possible (entanglement enhancement) [44, 39, 45, 46]. Previous control methods are mostly model-based: prior information about the system model is needed, such as the pulse method [34, 35, 36, 37, 38], time-continuous laser-driven approaches [39, 40], periodic modulations [41, 42, 43], optimal pulse protocols [44], linear quadratic-Gaussian (LQG) methods [39], and coherent feedback methods using auxiliary optical components [45, 46]. We note that there were two previous works [47, 48] on model-free RL for controlling and stabilizing a quantum system with an inverted harmonic potential and a double-well nonlinear potential, respectively, to a target state using weak-current measurements (WCMs) and partial state observation. However, these two works did not address entanglement control, while our work is developing a model-free deep-RL method to realize non-Gaussian entanglement engineering using only photon number counting from WCMs. (Backgrounds about WCM, deep RL, and quantum control are presented in Appendix A.) To our knowledge, prior to our work, model-free deep RL feedback control to create and stabilize the entanglement with WCM observations had not been available.
@@
II Results
II.1 Experimental proposal for entanglement engineering
@@
where $\hat{a}$ and $\hat{b}$ are the annihilation operators of the cavity and mechanical mode, respectively, $\hat{a}^{\dagger}$ and $\hat{b}^{\dagger}$ are the corresponding creation operators. The frequency detuning of the cavity is $\Delta\equiv\omega_{L}-\omega_{c}$. The nonlinear coupling $g_{0}$ arises from the radiation pressure force between the light and the movable mirror (details given in Appendix B), and $\alpha_{L}$ is the real amplitude of the driven electromagnetic field. We set $g_{0}>\kappa$ so that the single-photon optomechanical coupling rate $g_{0}$ exceeds the coupling strength $\kappa$ between the cavity and the vacuum bath. This condition guarantees observable nonlinear quantum effects [62]. Under the strong laser approximation: $|\bar{\alpha}_{c}|\gg 1$, where $|\bar{\alpha}_{c}|$ is the amplitude of the light field inside the cavity induced by the strong laser, we have $\hat{a}\approx\bar{\alpha}_{c}+\delta\hat{a}$ with $\delta\hat{a}$ denoting the excitation or the shifted oscillator on top of the large coherent state with the amplitude $\bar{\alpha}_{c}$. After the displacement transformation $e^{\bar{\alpha}_{c}\hat{a}^{\dagger}-\bar{\alpha}_{c}^{*}\hat{a}}\hat{a}e^{-\bar{\alpha}_{c}\hat{a}^{\dagger}+\bar{\alpha}_{c}^{*}\hat{a}}=\hat{a}-\bar{\alpha}_{c}\equiv\delta\hat{a}$, the resulting linearized beam-splitter or “swap” Hamiltonian  [63, 56] is
@@
The open dissipative quantum optomechanics under the WCM obey the stochastic master equation (SME) (see Sec. IV and Appendix C). The number $n_{\textnormal{traj}}$ of trajectories simulated from SME can be selected according to the following considerations. If the observable is some expected physical quantity, using one trajectory is sufficient to extract the information about the quantum state: $n_{\textnormal{traj}}=1$. Experimentally, WCMs are performed, encoding the Wiener process in the observation and resulting in a large variance from the expectation value. To reduce the variance, more quantum trajectories should be used. To make computations feasible, we use five trajectories: $n_{\textnormal{traj}}=5$.
@@
II.2 RL in linear quantum optomechanics
@@
with the associated expected photon number $\langle\hat{n}^{\textnormal{target}}_{p}\rangle=\langle\hat{P}_{1}\rangle=0.5$. Consequently, the reward function can be set as $R_{t}\equiv-|O_{t}-0.5|$, regardless of whether the observation $O_{t}$ is $\langle\hat{n}_{p}\rangle(t)$ or $\mathcal{I}(t)$. Because of the relatively small target value of the expected photon number: $\langle\hat{P}_{1}\rangle=0.5$, the variance $\mathcal{I}(t)$ in the WCM photocurrent can be reduced by a Gaussian filter [87] with the weak measurement rate $\eta\leq 1$. The Gaussian kernel parameters of the filter such as the filter interval and the variance can be numerically chosen to reduce the standard deviation of the measurement photocurrent into a certain range, e.g., about ten times larger than the mean value (details in Subsec. IV.2). The PPO agent applies an updated stochastic policy to the quantum optomechanical environment to maximize the accumulated reward, where the action $G(t)$ is proportional to the amplitude of the cavity mode: $G(t)=g_{0}\bar{\alpha}_{c}(t)$. The action can be controlled by an incident laser [64] and is continuous in a certain range, e.g., $G\in[-5,5]\omega_{m}$. The decay rate of cavity and mechanical modes $\kappa=0.01\,\omega_{m}$, $\gamma=0.01\,\kappa$, respectively, because the quality of the mechanical oscillator mode is generally better than that of the optical cavity or microwave resonator mode [62, 60, 59, 66, 67, 68].
@@
Our deep RL, a model-free learning method, is implemented in the measurement-based feedback control framework for entanglement engineering in open quantum optomechanics. Details about the PPO algorithm applied in the linear quantum optomechanics are presented in Appendix D. To appreciate its performance, we employ two benchmark methods for comparison: Bayesian [55, 48] and random control. Bayesian control [55, 48] is a state-based feedback control of the stochastic process as governed by the SME. In our case, the control law is given by $G(t)=-\lambda|\langle\hat{n}_{p}\rangle(t)-0.5|\omega_{m}$ with $\langle\hat{n}_{p}\rangle(t)$ being the observation, where the hyperparameter $\lambda$ can be numerically optimized based on the performance. If the observation is $\mathcal{I}(t)$, the control flow will be in the form $G(t)=-\lambda|\mathcal{I}(t)-0.5|\omega_{m}$, in which the Wiener process blocks the performance to some degree. In Bayesian control, the smaller the variance in the measured photocurrent, the better the performance. For the random control method, the flow is generated by a uniform distribution in the action range $G\in[-5,5]\,\omega_{m}$. Note that the actions $G$ of random control and deep RL are in the same range while the one of Bayesian control is determined by the hyperparameter $\lambda$ and the state-based observation value or the WCM photocurrent. To make a fair comparison, $\lambda_{opt}$ is optimized within the action range $G\in[-5,5]\omega_{m}$. Specifically, the optimized hyperparameter $\lambda_{opt}$ corresponds to the best performance of Bayesian control in the set $\lambda\in\{1,2,\ldots,\lambda_{max}\}$, where $\lambda_{max}$ is the maximum integer of $\lambda$ to guarantee the action range $G\in[-5,5]\,\omega_{m}$.
@@
II.3 RL in nonlinear quantum optomechanics
@@
II.4 Physical understanding of entanglement engineering through model-free deep RL
@@
II.4.1 Linear interaction regime
@@
II.4.2 Nonlinear interaction regime
@@
II.4.3 Weak continuous measurement
In an open quantum system, under WCM and quantum dissipation into the vacuum bath as well, a Wiener process occurs in the observable. More specifically, the Wiener process arises from the Gaussian-weighted projection over the eigenstates, which weakly extracts the partial information from the quantum system and induces stochastic disturbances in both the dynamical equation and observation. Such disturbances can avoid a complete quantum state collapse and provide the capability to extract the quantum information continuously in the time domain. However, the nonlinear stochastic process occurs in both quantum dynamical trajectories and the measurement photocurrent, making it challenging to control the quantum system through WCM continuously.
@@ III Discussion
Gate-based control uses quantum logic gates to manipulate qubit states and generate entanglement, typically relying on Von Neumann projective measurement, which requires $10^{3}-10^{6}$ repetitions to estimate an observable. In contrast, weak continuous measurement employs Gaussian projections, eliminating repeated system preparation and requiring only a single preparation for the entire time horizon. While adiabatic passage achieves high-fidelity states through slow evolution under a time-dependent Hamiltonian, its long duration remains a limitation. Advances in counterdiabatic driving [93] aim to address this limitation but remain model-based, relying on predefined mathematical models. Similarly, dissipation engineering [94], which leverages tailored dissipation to stabilize desired states, is also inherently model-based. The RL, by contrast, offers a model-free approach, requiring no prior knowledge of mathematical models and needing only a single initial preparation for each episode. Weak continuous measurement further reduces repetitions to one per time step, significantly lowering quantum resource use and running time, making the model-free framework efficient and adaptable across quantum models.
@@
In our work, all methods discussed operate within the framework of Markovian feedback, particularly RL based on Markov Decision Processes (MDPs). This means that the action at the next time step is determined only by the state at the current time step. However, the time delay is a fundamental challenge in any feedback control system, including quantum systems, arising from the inevitable lag between measurement, processing, and feedback application [95, 96, 97]. This delay often stems from the time required to extract information from the quantum system, such as through photodetectors or other measurement apparatus. To address this issue, incorporating delay-compensation techniques into the feedback controller design is essential to account for and mitigate the effects of known delays.
IV Methods
IV.1 Stochastic Master Equation
@@
where the Hamiltonian is $\tilde{H}=\tilde{H}_{bs}$ or $\tilde{H}_{nl}$ and $\rho$ is a density operator in the Hilbert space. Under the Born-Markov approximation [101, 102], which requires the system-bath coupling to be weak and the correlation time of the bath to be much shorter than a characteristic timescale of system-bath interactions, the Markovian master equation, i.e., the first two terms in the right-hand side of Eq. (IV.1), has the Lindblad form [101]. At absolute zero temperature, the following environmental operator $\mathcal{L}_{env}\;\rho$ can be introduced to describe the coupling between the system and vacuum bath: $\mathcal{L}_{env}\;\rho=\kappa\mathcal{D}(\hat{a})\rho+\gamma\mathcal{D}(\hat{b})\rho$, where the cavity and oscillator modes are coupled to the vacuum bath with the strength $\kappa$ and $\gamma$, respectively [57]. The deep RL results in the Lindblad master equation with the nonlinear interaction are presented in Appendix E.
The WCM process described by the last two terms in the right-hand side of Eq. (IV.1) is nonlinear and Markovian in the unconditional master equation [52] in $\rho$. Under WCM, a Wiener process $dW$ with a Gaussian distribution [52] arises from the Gaussian-weighted projection over the eigenstates that allows the quantum information to be extracted continuously in the time domain, subject to stochastic disturbances in the last term of Eq. (IV.1) and quantum decoherence in the penultimate term of Eq. (IV.1). (Appendix C provides a detailed derivation of the SME.) The Lindblad operator $\mathcal{D}$ and the measurement superoperator $\mathcal{H}$ in Eq. (IV.1) are given by
@@
IV.2 Implementation details of deep RL
For simulating the linear or nonlinear quantum optomechanical system described by Eq. (IV.1), we use the “taylor1.5” solver from the SME solver in the QuTip’s package [103] with the tolerance $\textnormal{tol}=10^{-6}$ and time step size $dt=0.01\,\omega_{m}^{-1}$. The measurement current is simulated with the “homodyne” method, and the custom environment is constructed by the open-source platform OpenAI-Gym [104]. For RL simulations, we construct the PPO agent [49] by “stable-baselines3” [105] in the A2C [106] settings, where stochastic policy (actor) and the value function (critic) are modeled by two independent neural network function approximators, i.e., a set of fully connected feed-forward networks of dimensions $256\times 128\times 64$ and the hyperbolic tangent nonlinear activation function for each hidden layer. For the nonlinear quantum optomechanical configuration, in the target-utilization phase, the recurrent PPO agent outperforms the PPO agent, where both independent critic and actor networks are MLPs followed by one independent layer of LSTM with $256\times 128\times 64$ fully connected networks and $256$ hidden states. More details, especially in the table of Gaussian kernel, are described in Appendix F.
Acknowledgments
We thank Dr. Kanu Sinha for discussions and comments. This work was supported by the Air Force Office of Scientific Research under Grant No. FA9550-21-1-0438 and by the Office of Naval Research under Grant No. N00014-24-1-2548. The Quantum Collaborative, led by Arizona State University, also provided valuable expertise and resources for this work through seed funding. The Quantum Collaborative connects top scientific programs, initiatives, and facilities with prominent industry partners to advance the science and engineering of quantum information science.
AUTHOR DECLARATIONS
Conflict of Interest
The authors have no conflicts to disclose.